# Quantifying and Optimizing Global Faithfulness in Persona-driven Role-playing

### v1

@@ 5.3 PRP as Simple Original Characters
Table 1: PRP Faithfulness Evaluation on simple and data contamination-free characters. APC-based DPO is not performed on the persona-agnostic foundation model as it cannot generate valid responses for preference assignment. CPO: Abbreviation of our Active-Passive-Constraint-based Direct Preference Optimization
@@ 5.4 PRP as Complicated Famous Figures
Table 2: PRP Faithfulness Evaluation ($\Delta$APC score) on characters with person statements at scale.
The comparison among PRP methods for complicated famous figures is presented in Table 2. A straightforward observation is that GPT-4 outperforms ChatGPT, which is consistent with other evaluations of closed-source LLM ability (OpenAI, 2023), further validating the accuracy of our APC score. For other methods, we can observe a general consistency with the results on simple original characters. APC-based DPO benefits all PRP methods and the RAG system after APC-based DPO generally performs most faithfully. EU leads to a performance drop since it encourages the model to stick to a single persona statement while the initial has the ability to combine multiple pieces of knowledge. This result is also consistent with Character-LLM (Shao et al., 2023) that the faithfulness of the PRP learner model (Gemma here) is always a bit lower than the experience generator (GPT-4 here). Since the PRP faithfulness gap narrows between open and closed-source LLMs, the effectiveness of EU also drops. Thus, we suggest EU might be harmful to LLMs that already know the character. Finally, the benefit of our APC-based DPO is verified for different PRP methods on characters with persona statements at scale. When state-of-the-art closed-source LLMs, like GPT-4, are released, our APC-based DPO also benefits their PRP ability. We continue the discussion on the full APC scores in Appendix F.
@@ 5.5 Property Analysis of PRP Methods
We perform APC-based analysis for PRP methods in Figure 3 to deepen our understanding of how different techniques affect PRP performance.
Scaling Rule of In-Context Personas
We first analyze how the scale of in-context persona information affects PRP faithfulness before or after APC-based DPO. We experiment on PRP as Eve for instance. The most effective in-context persona statement number is $5\sim 7$, and the PRP faithfulness drops with a longer context, showing the reason LCM is outperformed by RAG. Before APC-based DPO, a longer context ($8\sim 10$ persona statements) is even outperformed by very short contexts ($2\sim 3$ persona statements). After DPO, the faithfulness drop in longer contexts becomes less prominent, indicating the robustness improvement of LCM from APC-based DPO.
@@ 6 Case Study
While quantified results verify the advantages of our APC score criterion and APC-based DPO, performances on the practical PRP task need to be further reflected and discussed based on real cases. We include several cases here to cast deeper insight into how our APC benefits the PRP domain.
@@ 6.2 Violation Detection
We also use APC to specifically explain why EU sometimes uploads halluncinated information to PRP models. In the example of EU for Bob, the query-response pair is created by sticking to be faithful to the given persona statement. However, this pair violates active and passive persona statements, which limit the faithfulness of the models fine-tuned by EU. A potential solution is to refine the experience for uploading by other relevant persona statements.
@@ 7 Conclusion
This paper proposes a pioneering study on quantifying and optimizing the global faithfulness of PRP methods. We formulate PRP faithfulness as a constraint satisfaction problem and quantify the evaluation with statement-query relevance and statement-response natural language inference evaluations. Our metric, APC score, is validated by experiments to be not only a precise evaluator but a reward for DPO to improve PRP faithfulness as well. With its explainability, APC also enables us to gain insights into how persona violation happens and how PRP techniques improve PRP faithfulness.
Limitation and Future Work
@@
Acknowledgement
Appendix A Original Characters and Interview Queries
@@
Appendix B Symbolic Distillation
@@
We apply prompts in the Appendix G for symbolic distillation from GPT-4 to build statement-query relevance and statement-to-response NLI discriminators. The whole pipeline includes $4$ stages.
@@
These datasets, with statistics shown in Appendix E, are then used to fine-tune the discriminators. For the evaluation, the seed persona statements are based on three characters: Beethoven, Newton, and Socrates. For each character used to learn PRP methods, the datasets are prompted based on only the persona statements of that character. The RAG retriever is fine-tuned on the statement-query relevance dataset. For APC-based DPO, the discriminators are built in the same way as the evaluator. The hyperparameters are presented in Appendix C.
Appendix C More PRP Method Implementation Details
@@
Appendix D Human Evaluation
@@
Appendix E Statisitcs of Characters in Experiments
@@
Appendix F Full Award Result

### v2

@@ 5.3 PRP as Simple Original Characters
Table 1: PRP Faithfulness Evaluation on simple and data contamination-free characters. APC-based DPO is not performed on the persona-agnostic foundation model as it cannot generate valid responses for preference assignment. CPO: Abbreviation of our APC-based DPO
@@ 5.4 PRP as Complicated Famous Figures
Table 2: PRP Faithfulness Evaluation ($\Delta$APC score) on characters with persona statements at scale.
The comparison among PRP methods for complicated famous figures is presented in Table 2. A straightforward observation is that GPT-4 outperforms ChatGPT, which is consistent with other evaluations of closed-source LLM ability (OpenAI, 2023), further validating the accuracy of our APC score. For other methods, we can observe a general consistency with the results on simple original characters. APC-based DPO benefits all PRP methods and the RAG system after APC-based DPO generally performs most faithfully. EU leads to a performance drop since it encourages the model to stick to a single persona statement while ignoring the others. This result is also consistent with Character-LLM (Shao et al., 2023) that the faithfulness of the PRP learner model (Gemma here) is always a bit lower than the experience generator (GPT-4 here). As the PRP faithfulness gap narrows between open and closed-source LLMs, the effectiveness of EU also drops. Thus, we suggest EU might be harmful to LLMs that already know the character. Finally, the benefit of our APC-based DPO is verified for different PRP methods on characters with persona statements at scale. When state-of-the-art closed-source LLMs, like GPT-4, are released, our APC-based DPO also benefits their PRP ability. We continue the discussion on the full APC scores in Appendix G.
@@ 5.5 Property Analysis of PRP Methods
Scaling Rule of In-Context Persona Statements
As shown in Figure 3, we first analyze how the scale of in-context persona information affects PRP faithfulness before or after APC-based DPO. We experiment on PRP as Eve for instance. The most effective in-context persona statement number is $5\sim 7$, and faithfulness drops with a longer context, showing the reason LCM is outperformed by RAG. Before APC-based DPO, a longer context ($8\sim 10$ persona statements) is even outperformed by very short contexts ($2\sim 3$ persona statements). After DPO, faithfulness drops in longer contexts and becomes less prominent, indicating the robustness improvement of LCM from APC-based DPO.
@@ 6 Case Study
While quantified results verify the advantages of our APC score criterion and APC-based DPO, performances in practice have to be further reflected based on real cases. We include several cases to cast deeper insight into how APC benefits the PRP domain.
@@ 6.2 Violation Detection
We also use APC to specifically explain why EU sometimes uploads hallucinated information to PRP models. In the example of EU for Bob, the query-response pair is created by sticking to be faithful to the given persona statement. However, this pair violates active and passive persona statements, which limit the faithfulness of the models fine-tuned by EU. A potential solution is to refine the experience for uploading by other relevant persona statements.
@@ 7 Conclusion
This paper proposes a pioneering study on quantifying and optimizing the global faithfulness of PRP methods. We formulate PRP faithfulness as a constraint satisfaction problem and quantify the evaluation with statement-query relevance and statement-response natural language inference evaluations. Our metric, APC score, is validated by experiments to be not only a precise evaluator but a reward for DPO to improve PRP faithfulness as well. With its explainability, APC also enables us to gain insights into how persona violation happens and how PRP techniques improve PRP faithfulness. Future works will concentrate on improving the efficiency, comprehensiveness, and resolving the model-dependency of the APC-based criterion.
Acknowledgement
Appendix A Limitation and Future Work
@@
Appendix B Original Characters and Interview Queries
@@
Appendix C Symbolic Distillation
@@
We apply prompts in the Appendix H for symbolic distillation from GPT-4 to build statement-query relevance and statement-to-response NLI discriminators. The whole pipeline includes $4$ stages.
@@
These datasets, with statistics shown in Appendix F, are then used to fine-tune the discriminators. For the evaluation, the seed persona statements are based on three characters: Beethoven, Newton, and Socrates. For each character used to learn PRP methods, the datasets are prompted based on only the persona statements of that character. The RAG retriever is fine-tuned on the statement-query relevance dataset. For APC-based DPO, the discriminators are built in the same way as the evaluator. The hyperparameters are presented in Appendix D.
Appendix D More PRP Method Implementation Details
@@
Appendix E Human Evaluation
@@
Appendix F Statisitcs of Characters in Experiments
@@
Appendix G Full Award Result
@@ Appendix H Prompts
Appendix I More Characters
Besides characters in the main content, we further expand the scope of characters to different ethnicity, which involves,
  • Alex: An African American baseball player
  • Isabella: An Italian traveling cook
  • Takayoshi: A Japanese game developer
  • Ousmane: A rich gold mine owner of the Malian Empire in the 1300s
  • Jones: A young British worker in the Victorian Era
  • Zhe: A Chinese poet in the Tang Dynasty
  • Crossan: A time-traveling scientist
  • Betty: A pet cat who can talk with ghosts
  • X: An alien space traveler and photographer
These characters can better represent people with different spatial and temporal backgrounds and even cover non-human characters from the fantasy world.
Character Alex Isabella Takayoshi Ousmane Jones Zhe Crossan Betty X
w/o CPO Vanilla 0.5 0.8 0.6 0.3 0.9 1.1 0.3 0.3 0.7
EU 1.8 2.8 2.0 1.4 0.7 3.8 2.0 1.2 5.2
LCM 7.1 7.4 6.5 4.5 6.2 5.2 2.2 2.8 8.1
RAG 7.6 8.1 6.9 3.0 6.6 5.8 1.8 3.2 7.5
w/ CPO EU 5.3 6.1 5.7 3.6 4.8 4.9 3.1 2.9 7.9
LCM 7.5 7.7 7.0 4.8 6.2 5.4 4.5 3.9 8.2
RAG 7.9 8.2 7.4 3.9 7.5 6.9 2.5 4.6 8.9
Table 6: PRP performance on more characters based on the distilled DeBERTa Evaluator
Character Alex Isabella Takayoshi Ousmane Jones Zhe Crossan Betty X
w/o CPO Vanilla 0.2 0.1 -0.2 0.5 0.2 0.2 0.4 0.1 0.8
EU 1.4 1.8 3.0 0.5 1.3 6.4 1.2 0.3 7.4
LCM 3.1 8.6 5.6 4.1 7.4 3.4 2.1 1.6 11.3
RAG 3.3 7.8 6.1 1.6 8.1 4.3 2.7 2.2 10.1
w/ CPO EU 2.7 5.6 5.9 3.0 4.7 7.1 2.2 1.5 9.5
LCM 3.2 9.8 8.1 4.6 8.2 7.8 4.0 2.3 12.1
RAG 4.8 10.0 9.8 2.0 8.3 7.3 2.9 3.1 14.6
Table 7: PRP performance on more characters based on the GPT-4 Evaluator
The experiment results are presented in Tables 6 and 7, which is consistent with our results in Tables 1 and 2. Thus, our conclusion is certificated on a larger scope for broader application.
Appendix J Metric Comparison
To better justify selecting our APC score and also support the claim that the fine-grained APC score has the advantage over coarse-grained metrics, we add a coarse-grained metric as the baseline. We directly prompt GPT-4 with the criterion used for human evaluation shown in Appendix E. We also distill this scoring ability (following the same scenario as APC) to DeBERTa to check whether the efficiency can be boosted. We evaluate the Spearman correlation between the metric and the human evaluation of the 7 role-playing methods on the 3 human-evaluated characters.
Character (#Persona Statement) Alice (8) Bob (19) Eve (30)
GPT-4 Coarse-grained Score 92.42 86.27 81.40
APC Score 97.18 99.10 99.10
DeBERTa Coarse-grained Score 81.40 69.91 54.57
APC Score 88.61 95.50 99.10
Table 8: Comparison of PRP metrics on the consistency with human evaluation.
The results verify that 1) Fine-grained APC score shows better consistency with human evaluation. 2) The fine-grained APC score is stable to the number of persona statements while the coarse-grained score degrades with the increase of persona statements. 3) The coarse-grained evaluating ability is harder to be distilled into smaller models for efficiency boosting. Based on case checking, we find an underlying issue of the coarse-grained metric is the LLM will assign a high score to a response once it contains some correct information, ignoring the missing important information (active constraint) and occasionally conflictions (passive constraint).
Appendix K Student Model Comparison
We select DeBERTa as the student model to distill from GPT-4 because small encoders (BERT, RoBERTa, etc.) show promising performance on relevance and NLI, which are classic NLU tasks in the GLUE benchmark. Among encoders, DeBERTa (DeBERTa-v3-large) is a state-of-the-art model that shows strong performance after fine-tuned on NLU tasks. To further verify DeBERTa as a proficient student model, we add an analysis of the in-domain (ID)/out-of-domain (OOD) performance and the efficiency of different base models for distillation.
Task Relevance NLI Efficiency
ID OOD ID OOD
DeBERTa (Base) 92.46 89.90 89.72 87.80 409.6it/s
DeBERTa (Large) 94.04 92.10 93.46 91.50 150.8it/s
Gemma-1.1-it (2b) 94.25 92.50 93.68 91.80 26.4it/s
Table 9: Model Performance Comparison
The in-domain test set (1697 instances for Relevance, 3773 instances for NLI) is the 20% split of the characters (Beethoven, Newton, Socrates) that build the training set (6787 instances for Relevance, 15092 instances for NLI). The out-of-domain test set samples 1000 cases from other characters. The results show DeBERTa-V3-Large (300M) shows a comparative performance with a 2B Gemma model, while is about 6 times faster, which justifies DeBERTa to be a strong student model. The out-of-domain performance is generally high, which indicates the generalizability to other characters. Finally, an extra discovery is that DeBERTa-v3-base (100M) can further significantly boost efficiency with some trade-offs in accuracy.
NeurIPS Paper Checklist
The checklist is designed to encourage best practices for responsible machine learning research, addressing issues of reproducibility, transparency, research ethics, and societal impact. Do not remove the checklist: The papers not including the checklist will be desk rejected. The checklist should follow the references and follow the (optional) supplemental material. The checklist does NOT count towards the page limit.
Please read the checklist guidelines carefully for information on how to answer these questions. For each question in the checklist:
  • You should answer [Yes] , [No] , or [N/A] .
  • [N/A] means either that the question is Not Applicable for that particular paper or the relevant information is Not Available.
  • Please provide a short (1–2 sentence) justification right after your answer (even for NA).
The checklist answers are an integral part of your paper submission. They are visible to the reviewers, area chairs, senior area chairs, and ethics reviewers. You will be asked to also include it (after eventual revisions) with the final version of your paper, and its final version will be published with the paper.
The reviewers of your paper will be asked to use the checklist as one of the factors in their evaluation. While "[Yes] " is generally preferable to "[No] ", it is perfectly acceptable to answer "[No] " provided a proper justification is given (e.g., "error bars are not reported because it would be too computationally expensive" or "we were unable to find the license for the dataset we used"). In general, answering "[No] " or "[N/A] " is not grounds for rejection. While the questions are phrased in a binary way, we acknowledge that the true answer is often more nuanced, so please just use your best judgment and write a justification to elaborate. All supporting evidence can appear either in the main paper or the supplemental material, provided in appendix. If you answer [Yes] to a question, in the justification please point to the section(s) where related material for the question can be found.
IMPORTANT, please:
  • Delete this instruction block, but keep the section heading “NeurIPS paper checklist",
  • Keep the checklist subsection headings, questions/answers and guidelines below.
  • Do not modify the questions and only use the provided macros for your answers.
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We clearly show the claims in the abstract and introduction, which is explored and verified in experiments and analyses.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: You can refer to the limitation section.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: We do not include theoretical results.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We include all hyperparameters and other settings for the reproduction of our results.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We use open-source tools to implement the experiments, with clear instructions for reproduction.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We include all hyperparameters and other settings for the reproduction of our results.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: The improvement shown in our experiments is statistically significant.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We mention the devices used for computer resources.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We follow the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We discuss the broader impacts of the work.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [Yes]
  54. Justification: We discuss the safeguard of the work.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We discuss the licenses for existing assets.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We document the new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [Yes]
  69. Justification: We present a clear guideline for human evaluation.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.